OpenStudy (anonymous):

Which is correct? 1. sqrt(-4) * sqrt(-16) = sqrt(-16*-4) = sqrt(64) = 8. OR 2. sqrt(-4) * sqrt(-16) = 2i * 4i = 8i^2 = -8?

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

why, is there something wrong with the first one?

OpenStudy (anonymous):

only time \[\sqrt{a}\sqrt{b}=\sqrt{ab}\] is if a and b are non-negative

OpenStudy (anonymous):

\[\sqrt{-4}=2i\] \[\sqrt{-16}=4i\] \[2i\times 4i=8i^2=-8\]

OpenStudy (anonymous):

wow I never knew that that the rule was for non-negative numbers only. But why is it for non-negative numbers only?

OpenStudy (anonymous):

I mean if I didnt know about complex numbers then would the first one be correct?